# Prominence-Aware Artifact Detection and Dataset for Image Super-Resolution

### v1

@@ Appendix E Subjective evaluation on additional SR datasets
We conduct an additional subjective evaluation on 6 widely known image datasets (Martin et al., 2001; Wang et al., 2018; Dong & Loy, 2016; Bevilacqua et al., 2012; Yang et al., 2010), following the setup described in Section 5.3. In total, this evaluation used 420 source images, each processed by 8 SR models.
Tables 11 and 11 show the results, grouped by SR models and by artifact detection methods, respectively. Interestingly, SR models show much better artifact robustness than in our main comparison in Section 5.3, likely because these datasets are commonly used for SR training and evaluation. Our proposed method falls one confident artifact short of DISTS, but otherwise shows competitive results.
Appendix F F1-score bounds analysis
@@
Appendix G Objective evaluation on the learning-based compression task
@@
In order to evaluate the transferability of our promised method to other domains, we conduct an objective evaluation on the JPEG AI edge artifact examples dataset collected by Tsereh et al., following our methodology described in Section 5.1. Tables 13 and 13 show the results.
@@
Appendix H Additional experiments
@@
This appendix describes our additional experiments and ablation studies. For all experiments, we measure objective metrics as we described in Section 5.1 of the main paper. Tables 14 and 16 show the results on our proposed dataset, and Tables 15 and 17 show the results on the DeSRA (Xie et al., 2023a) dataset.
@@
H.1 Input-feature ablation
@@
H.2 Training on different pseudo-GT
@@
H.3 GT-area-only training
@@
H.4 Weighting loss to normalize training-set classes
@@
H.5 Variance across training runs
@@
H.6 Using a CNN instead of a per-pixel MLP
@@
H.7 Using a random forest instead of MLP
@@
H.8 Filtering heatmaps through semantic segmentation
Similarly to Xie et al. (2023a), we noticed that SR-artifact prominence is related to the affected object. Viewers hardly notice artifacts in objects such as grass, leaves, dirt, and soil. To confirm this hypothesis, we used the SAN (Xu et al., 2023) semantic segmentation method to cancel artifact predictions on pixels of certain semantic classes.

### v2

@@ Appendix F Subjective evaluation on additional SR datasets
We conduct an additional subjective evaluation on 6 widely known image datasets (Martin et al., 2001; Wang et al., 2018; Dong and Loy, 2016; Bevilacqua et al., 2012; Yang et al., 2010), following the setup described in Section 5.4. In total, this evaluation used 420 source images, each processed by 8 SR models.
Tables 12 and 12 show the results, grouped by SR models and by artifact detection methods, respectively. Interestingly, SR models show much better artifact robustness than in our main comparison in Section 5.4, likely because these datasets are commonly used for SR training and evaluation. Our proposed method falls one confident artifact short of DISTS, but otherwise shows competitive results.
Appendix G F1-score bounds analysis
@@
Appendix H Objective evaluation on the learning-based compression task
@@
In order to evaluate the transferability of our promised method to other domains, we conduct an objective evaluation on the JPEG AI edge artifact examples dataset collected by Tsereh et al., following our methodology described in Section 5.1. Tables 14 and 14 show the results.
@@
Appendix I Additional experiments
@@
This appendix describes our additional experiments and ablation studies. For all experiments, we measure objective metrics as we described in Section 5.1 of the main paper. Tables 15 and 17 show the results on our proposed dataset, and Tables 16 and 18 show the results on the DeSRA (Xie et al., 2023) dataset.
@@
I.1 Input-feature ablation
@@
I.2 Training on different pseudo-GT
@@
I.3 GT-area-only training
@@
I.4 Weighting loss to normalize training-set classes
@@
I.5 Variance across training runs
@@
I.6 Using a CNN instead of a per-pixel MLP
@@
I.7 Using a random forest instead of MLP
@@
I.8 Filtering heatmaps through semantic segmentation
Similarly to Xie et al. (2023), we noticed that SR-artifact prominence is related to the affected object. Viewers hardly notice artifacts in objects such as grass, leaves, dirt, and soil. To confirm this hypothesis, we used the SAN (Xu et al., 2023) semantic segmentation method to cancel artifact predictions on pixels of certain semantic classes.